OpenStudy (anonymous):

Please Help!! Determine whether the sequence converges or diverges. If it converges, give the limit. 48, 8, 4/3, 2/9

OpenStudy (anonymous):

@SolomonZelman @precal @paki

OpenStudy (jdoe0001):

hmm what does it mean to converge?

OpenStudy (anonymous):

It means that the sum of the individual terms of the sequence from n=0 to infinity equates to some finite number.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

yes... so .... notice the numbers.... from 48 to 8 is a division by 6 from 8 to 4/3 is a division by 6 from 4/3 to 2/9 is a division by 6 so we could say that \(\bf a_1 = 48\quad r=6\) thus \(\Large \bf S=\Sigma_{i=0}^\infty\ a_1{\color{brown}{ r}}^i\implies \cfrac{a_1}{1-{\color{brown}{ r}}}\implies \cfrac{48}{1-{\color{brown}{ 6}}}\)

OpenStudy (jdoe0001):

and that reaches a limit, a finite value, thus it "converges"

OpenStudy (jdoe0001):

hmm actually.... wait.. a sec... I may have spoken too soon our "r" should be 1/6 not 6 lemme redo that

OpenStudy (jdoe0001):

1/6 since is a "division" by 6

OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 48, 8, 4/3, 2/9, ... Converges; 288/5 Converges; 0 Diverges Converges; -12432

OpenStudy (anonymous):

those are my only choices..

OpenStudy (anonymous):

oh ok

OpenStudy (jdoe0001):

\(\large \bf S=\Sigma_{i=0}^\infty\ a_1{\color{brown}{ r}}^i\implies \cfrac{a_1}{1-{\color{brown}{ \frac{1}{6}}}}\implies \cfrac{48}{1-{\color{brown}{ \frac{1}{6}}}}\implies \cfrac{48}{\frac{5}{6}}\implies \cfrac{48}{1}\cdot \cfrac{6}{5}\)

OpenStudy (jdoe0001):

hmm got a bit truncated so \(\bf \large { S=\Sigma_{i=0}^\infty\ a_1{\color{brown}{ r}}^i\implies \cfrac{a_1}{1-{\color{brown}{ \frac{1}{6}}}}\implies \cfrac{48}{1-{\color{brown}{ \frac{1}{6}}}}\implies \cfrac{48}{\frac{5}{6}} \\ \quad \\ \implies \cfrac{48}{1}\cdot \cfrac{6}{5} }\)

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

so its 288/5 as the limit

OpenStudy (jdoe0001):

yeap... it has a finite end. and thus it converges, it converges at 288/5

OpenStudy (anonymous):

thank you :)